Can someone please walk me through this?
Consider the table below.
x f(x) x g(x)
2 -8 -9 2
5 4 -8 -4
Use the table of values for functions f and g to find the composition (f compose g)(-9).
Jerry
jerry@san.rr.com

Found 2 solutions by Fombitz, jessica43:
Answer by Fombitz(32388)
Answer by jessica43(140)

What is "(f o g )(x)"? This is read as "f-compose-g of x", and means "plug x into g, evaluate, and then plug the result into f". The computation can feel a lot easier if I use the following, more intuitive, formatting:
(f o g )(x) = f( g(x))
So in this problem, we have (f o g )(-9). This means the first step is to plug in -9 into the g(x) function.
Using the given table, we see that when x=-9, then g(x) = 2. So g(-9) = 2.
The next step is to take this answer (2), and plug it into f(x).
Again using the table, since x=2, then f(x)=-8.
So, (f o g )(-9) = f(g(-9)) = f(2) = -8.
Your answer is -8.